OpenStudy (bluescar):

Enter an equation that can be used to solve this problem, where x represents the number of hours that the first plane had been traveling. Two planes left the same airport traveling in opposite directions. The first plane left the airport at 2 P.M. traveling at an average rate of 500 mph. The second plane left the airport at 4 P.M. traveling at an average rate of 550 mph. When will the planes be 3100 miles apart? Enter the amount hours, after the first plane left, when the planes will be 3100 miles apart. Equation: 2 h

OpenStudy (phi):

do you know about rate * time = distance ?

OpenStudy (bluescar):

no XD

OpenStudy (phi):

if you go 500 mph for 2 hours, any idea how far you went?

OpenStudy (bluescar):

1000 miles

OpenStudy (phi):

yes, you did 500*2 if you go for "x" hours you do the same thing: 500*x

OpenStudy (bluescar):

oh

OpenStudy (phi):

in other words, The first plane left the airport at 2 P.M. traveling at an average rate of 500 mph for "x" hours we don't know what x is (as a number) but we know the distance will be 500x ok so far ?

OpenStudy (bluescar):

mhm

OpenStudy (phi):

distance is rate * time rate (or speed) is 500 miles per hour time is x (unknown) number of hours the distance is 500*x or 500x (we usually leave out the multiply sign, but 500x means 500 * x)

OpenStudy (bluescar):

I'm still with you not confused yet XD

OpenStudy (phi):

the 2nd plane leaves two hours later if the first plane flies for x hours, any idea how long the 2nd plane is flying ?

OpenStudy (bluescar):

550 is how fast the next plane is flying

OpenStudy (phi):

yes, but let's say the first plane flies for x hours and then we check how long the 2nd plane has been in the air (assuming it's still flying) any idea how long the 2nd plane flew ?

OpenStudy (phi):

let's use numbers if the first plane has been flying for 3 hours, at that same time, how long has the 2nd plane been in the air?

OpenStudy (bluescar):

it flew 1650

OpenStudy (phi):

ok, but that is not the question. if the first plane leaves at 2 pm and flies for 3 hours (so it's now 5 pm) and the 2nd plane leaves at 4 pm, how long has the 2nd plane been flying ?

OpenStudy (bluescar):

1h

OpenStudy (phi):

if the first plane flies for 5 hours how long will the 2nd plane be flying ? any idea ?

OpenStudy (bluescar):

2500

OpenStudy (phi):

how many hours ?

OpenStudy (bluescar):

5

OpenStudy (phi):

5 hours for the first plane how many hours for the 2nd plane ?

OpenStudy (bluescar):

3

OpenStudy (phi):

yes. how did you get 3?

OpenStudy (phi):

you did 5 - 2

OpenStudy (bluescar):

mhm

OpenStudy (phi):

the idea is the 2nd plane flies 2 less hours because it leaves 2 hours later

OpenStudy (bluescar):

so the second plan flies 1650

OpenStudy (phi):

ok, but now we want to know how long it flies if the first plane flies for "x" hours we use the idea that the 2nd plane flies 2 less hours or (x-2) hours.

OpenStudy (phi):

next, we want the distance that the 2nd plane flies if it flies for (x-2) hours how do we figure that out? we multiply the time (x-2) times the speed 550 can you do that ?

OpenStudy (bluescar):

so we need to keep multipling to the second plane to figure it out ?

OpenStudy (phi):

we use rate * time to find the distance rate (or speed) is 550 time is (x-2)

OpenStudy (phi):

in other words, the distance the 2nd plane goes is 550(x-2)

OpenStudy (bluescar):

mhm

OpenStudy (phi):

the say the planes go in opposite directions, so the distance between them is the sum of how far each plane flies distance = 500x + 550(x-2)

OpenStudy (phi):

they want to know how many hours to get a distance of 3100 miles 3100 = 500x + 550(x-2) that is the equation

OpenStudy (bluescar):

i think i understand and the time is 3h

OpenStudy (phi):

if we use x=3 in the formula 500x + 550(x-2) we would get 500*3 + 550(3-2) if we simplify that we get 1500+ 550*1 or 2050 that is not far enough

OpenStudy (bluescar):

oh ok thanks !

OpenStudy (phi):

we can use algebra to "solve for x" in the formula 500x + 550(x-2) = 3100 I would first "distribute" the 550 and do 550*x and 550*-2 500x + 550x -1100= 3100 next add 500 x's plus 550 more x's. that is 1050 x's or 1050x 1050x -1100 = 3100 now add +1100 to both sides.

OpenStudy (bluescar):

2150

OpenStudy (phi):

you should get 1050x = 3100+1100 or 1050x = 4200 finally, divide both sides by 1050

OpenStudy (bluescar):

1050 for the left side gives you 1 and on the right gives 4

OpenStudy (phi):

almost. 1050x/1050 is 1x or x don't ignore the letters. x= 4 is the answer it takes 4 hours for the two planes to get 3100 miles apart.

OpenStudy (bluescar):

ahh

OpenStudy (bluescar):

thank you and bye XD